OpenStudy (anonymous):

Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2? helpp !!!

OpenStudy (anonymous):

y=kx where k is the proportionality constant so 8=-6k so,k=-4/3 so, when x=-2 y=-4/3*-2 =8/3

OpenStudy (anonymous):

would this b it The slope is –2 and the y-intercept is –4.

OpenStudy (anonymous):

Yeaa ,

OpenStudy (anonymous):

^_^ Thank you :P

OpenStudy (anonymous):

Youu Welcome ")